# Preserving Expert-Level Privacy in Offline Reinforcement Learning

### v1

@@ B.2 Expert-Level DP-SGD
1:Model Parameters $\phi$, Expert Datasets $D^{\Pi}$, Expert Policies $\Pi$, Loss function $\mathcal{L(\phi)}=\frac{1}{b}\sum_{i=1}^{b}\mathcal{L}(\phi,x_{i})$, Learning Rate $\eta$, Batch size $b$, Noise parameter $\sigma$, Clipping threshold $C$, Training steps $N$
@@
In our method, Algorithm 4 is incorporated in the training scheme in Algorithm 3 (lines 5 and 6) in the main paper. At each training step, with probability $p$, we execute lines 3-8 in Algorithm 4 to train on a batch of transitions. Otherwise, we sample a batch of transitions from $\mathcal{D}_{stable}^{\Pi}$ and perform a normal SGD update (this has no privacy cost). The privacy analysis for Algorithm 3 proceeds in a manner similar to that in standard DP-SGD (Abadi et al.,, 2016) with the exception of the subsampling probability which must be modified to be $q=p\cdot\frac{b}{m}$ (in the analysis only) to account for the randomness in picking the unstable dataset for sampling.
@@ B.3 Hyperparameter Tuning
For all the environments, we assume that $p_{min}=0.02$. For LunarLander and CartPole, for $\varepsilon\leq 5$, we set $\varepsilon_{1}=0,\varepsilon_{2}=\varepsilon,\delta_{1}=0,\delta_{2}=\delta$ and resort to pure expert-level DP-SGD training since the privacy budget is not enough to get meaningful stable trajectories for training. Similarly for Acrobot, for $\varepsilon\leq 2.5$, we omit stable data collection and just do expert-level DP-SGD training. For HIV Treatment and for all other values of $\varepsilon$ for Lunar Lander and CartPole, we set $\varepsilon_{1}=\frac{3\varepsilon}{4},\varepsilon_{2}=\frac{\varepsilon}{4},\delta_{1}=\frac{9\delta}{10},\delta_{2}=\frac{\delta}{10}$. For Acrobot, we set $\varepsilon_{1}=\varepsilon,\varepsilon_{2}=0,\delta_{1}=\delta,\delta_{2}=0$.
@@
We use Conservative Q-Learning (CQL) (Kumar et al., (2020)) for training on the offline data. For Lunar Lander, Acrobot and CartPole, we fix the batch size to 256, and the neural network hidden layer sizes to (256, 256). For HIV Treatment, we set the batch size as 64 and use a neural network of one hidden layer of size 512 to represent the Q-function. For DP-SGD updates, we clip all gradient norms to 1.0. We tune the following 3 hyperparameters for our method: learning rate $\eta$, probability of sampling from $\mathcal{D}_{unst}^{\Pi}$ during training ($p$), and DP-SGD noise multiplier which is the standard deviation of the gaussian noise applied to the gradients divided by the clipping threshold. We perform a grid search over all hyper-parameters. The search spaces for different hyperparameters are as follows: $\eta$ [0.0001, 0.001, 0.005, 0.01, 0.05], $p$ [0.9, 0.8, 0.5, 0.0], and noise multiplier - [10.0, 20.0, 30.0, 50.0, 60.0 80.0]. For each configuration of hyperparameters, we let the model train for as long as possible with the given value of $\varepsilon_{2}$. We report the average episodic return obtained over 10 evaluation runs spaced over the last 10000 steps of training.
Acrobot: We fix $p=0.0$, that is, we use only $\mathcal{D}_{stable}^{\Pi}$ for training. Over all values of $\varepsilon$, (5, 7.5, 10), $\eta$ = 0.005 gave the best performance. Since $p=0.0$, the noise multiplier is not required. For $\varepsilon=$ 10, 7.5, 5, we set $T$ (defined in Algorithm 2) equal to 25, 20, 20 respectively.
Lunar Lander: For $\varepsilon=10.0$, the best performing values of $\eta$, $p$ and noise multiplier are 0.005, 0.8, 60.0 respectively. For $\varepsilon=7.5$, the best performing hyperparameters, in the same order, are 0.005, 0.5, 80.0. For $\varepsilon=5.0,2.5$, we do pure expert-level DP-SGD training. For $\varepsilon=$ 10, 7.5, we set $T$ equal to 100, 62 respectively.
CartPole: For $\varepsilon=10.0$, the best performing values of $\eta$, $p$ and noise multiplier are 0.001, 0.9, 50.0 respectively. For $\varepsilon=7.5$, the best performing hyperparameters, in the same order, are 0.001, 0.8, 60.0. For lower values of $\varepsilon$, we do pure expert-level DP-SGD training. For $\varepsilon=$ 10, 7.5, we set $T$ equal to 25, 20 respectively.
HIV Treatment: We fix $p=0.0$, that is, we use only $\mathcal{D}_{stable}^{\Pi}$ for training (hence noise multiplier is not required). For $\varepsilon=10.0$, learning rate = 0.0001 gave the best performance. We use $T=50$.
For tuning the expert-level DP-SGD baseline, we vary $\eta$ and noise multiplier over the same ranges as specified above. We again let the model train for as long as possible with the given value of $\varepsilon$, and report the average episodic return obtained over 10 evaluation runs spaced over the last 10000 steps of training.

### v2

@@ B.2 Expert-Level DP-SGD
1:Model Parameters $\phi$, Expert Datasets $D^{\Pi}$, Expert Policies $\Pi$, Loss function $\mathcal{L(\phi)}=\frac{1}{b}\sum_{i=1}^{b}\mathcal{L}(\phi,x_{i})$, Learning Rate $\eta$, Batch size $b$, Noise parameter $\sigma$, Clipping threshold $C$, Training steps NN
@@
In our method, Algorithm 4 is incorporated in the training scheme in Algorithm 3 (lines 5 and 6) in the main paper. At each training step, with probability $p$, we execute lines 3-8 in Algorithm 4 to train on a batch of transitions. Otherwise, we sample a batch of transitions from $\mathcal{D}_{stable}^{\Pi}$ and perform a normal SGD update (this has no privacy cost). The privacy analysis for Algorithm 3 proceeds in a manner similar to that in standard DP-SGD (abadi2016dp) with the exception of the subsampling probability which must be modified to be $q=p\cdot\frac{b}{m}$ (in the analysis only) to account for the randomness in picking the unstable dataset for sampling.
@@ B.3 Hyperparameter Tuning
For all the environments, we assume that $p_{min}=0.02$. For LunarLander, CartPole and HIVTreatment, for $\varepsilon=5$, we set $\varepsilon_{1}=0,\varepsilon_{2}=\varepsilon,\delta_{1}=0,\delta_{2}=\delta$ and resort to pure expert-level DP-SGD training since the privacy budget is not enough to get meaningful stable trajectories for training. For Acrobot, we set $\varepsilon_{1}=\varepsilon,\varepsilon_{2}=0,\delta_{1}=\delta,\delta_{2}=0$. For all other cases, we set $\varepsilon_{1}=\frac{3\varepsilon}{4},\varepsilon_{2}=\frac{\varepsilon}{4},\delta_{1}=\frac{9\delta}{10},\delta_{2}=\frac{\delta}{10}$.
We use Conservative Q-Learning (CQL) (kumar2020conservative) and Discrete Batch-Constrained Deep Q-Learning (BCQ) (fujimoto2019benchmarking) for training on the offline data. For Lunar Lander, Acrobot and CartPole, we fix all neural network hidden layer sizes to (256, 256). For HIV Treatment, we use a neural network of one hidden layer of size 512 to represent the Q-function. For DP-SGD updates, we clip all gradient norms to 1.0. For Acrobot, for $\varepsilon=$ 10, 7.5, 5, we set $T$ (defined in Algorithm 2) equal to 25, 20, 20 respectively. For LunarLander, for $\varepsilon=$ 10, 7.5, we set $T$ equal to 100 and 62 respectively. For CartPole, for $\varepsilon=$ 10, 7.5, we set $T$ equal to 25 and 20 respectively. We use $T=50$ for HIVTreatment.
@@
We tune the following hyperparameters for our method: learning rate $\eta$, batch size $b$, probability of sampling from $\mathcal{D}_{unst}^{\Pi}$ during training ($p$), and DP-SGD noise multiplier ($s$) which is the standard deviation of the gaussian noise applied to the gradients divided by the clipping threshold. We perform a grid search over all hyper-parameters. The search spaces for different hyperparameters are as follows: $\eta$ - [0.0001, 0.0005, 0.001, 0.005, 0.01], $b$ - [64, 128, 256], $p$ - [0.9, 0.8, 0.5, 0.0], and $s$ - [10.0, 20.0, 30.0, 40.0, 50.0, 80.0]. For each configuration of hyperparameters, we let the model train for as long as possible with the given value of $\varepsilon_{2}$. We report the average episodic return obtained over 10 evaluation runs spaced over the last 10000 steps of training. The best hyper-parameter configurations for each setting are given in Table 3.
Table 3: Best hyper-parameter settings across different environments, offline-RL algorithms and $\varepsilon$ values.
Environment $\mathbf{\varepsilon}$ CQL BCQ
$\eta$ $b$ $p$ $s$ $\eta$ $b$ $p$ $s$
Acrobot 5 0.001 128 0.0 - 0.01 256 0.0 -
7.5 0.001 128 0.0 - 0.01 256 0.0 -
10 0.0005 128 0.0 - 0.01 256 0.0 -
LunarLander 5 0.0005 256 1.0 80 0.005 64 1.0 10
7.5 0.005 64 0.8 20 0.001 128 0.5 50
10 0.0005 64 0.8 40 0.005 128 0.5 20
CartPole 5 0.005 256 1.0 20 0.005 128 1.0 10
7.5 0.001 128 0.9 50 0.01 128 0.5 20
10 0.0001 128 0.8 80 0.001 128 0.5 40
HIVTreatment 5.0 0.01 256 1.0 10 0.005 128 1.0 20
7.5 0.0001 64 0.0 - 0.005 64 0.9 20
10 0.0001 64 0.0 - 0.005 64 0.9 20